Metlock Inc. has issued three types of debt on January 1, 2020, the start of the company’s fiscal year.

(a) $10 million, 10-year, 13% unsecured bonds, interest payable quarterly. Bonds were priced to yield 11%.
(b) $27 million par of 10-year, zero-coupon bonds at a price to yield 11% per year.
(c) $15 million, 10-year, 10% mortgage bonds, interest payable annually to yield 11%.


Prepare a schedule that identifies the following items for each bond: (1) maturity value, (2) number of interest periods over life of bond, (3) stated rate per each interest period, (4) effective-interest rate per each interest period, (5) payment amount per period, and (6) present value of bonds at date of issue. (Round stated and effective rate per period to 2 decimal places, e.g. 10.25%. Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971.)

Unsecured
Bonds

Zero-Coupon
Bonds

Mortgage
Bonds

(1) Maturity value $ $ $
(2) Number of interest periods
(3) Stated rate per period % % %
(4) Effective rate per period % % %
(5) Payment amount per period $ $ $
(6) Present value $ $ $
Answer #1

Unsecured
Bonds

Zero-Coupon
Bonds

Mortgage
Bonds

(1)

Maturity value

$10000000

$27000000

$15000000

(2)

Number of interest periods

40 (10*4)

10

10

(3)

Stated rate per period

3.25% (13%/4)

0%

10%

(4)

Effective rate per period

2.75% (11%/4)

11%

11%

(5)

Payment amount per period

$325000

0

$1500000

(6)

Present value

$11203883

$9508860

$14116545

Payment amount per period

Unsecured Bonds = 10000000*3.25% = 325000

Zero-Coupon Bonds = 27000000*0% = 0

Mortgage Bonds =15000000*10% = 1500000

Present value

Unsecured Bonds

Present Value of Face Value = 10000000*0.33785 = 3378500

Present value of interest payment = 325000*24.07810 = $7825383

Present value = 3378500+7825383 = 11203883

Present value factor of $1 @ n= 40, I = 2.75% is 0.33785

Present value ordinary annuity of $1 @ n= 40, I = 2.75% is 24.07810

Zero-Coupon Bonds

Present Value of Face Value = 27000000*0.35218 = 9508860

Present value factor of $1 @ n= 10, I = 11% is 0.35218

Mortgage Bonds

Present Value of Face Value = 15000000*0.35218 = 5282700

Present value of interest payment = 1500000*5.88923 = $8833845

Present value = 5282700+8833845 = 14116545

Present value factor of $1 @ n= 10, I = 11% is 0.35218

Present value ordinary annuity of $1 @ n= 10, I = 11% is 5.88923